OpenStudy (anonymous):

Quick question. One number exceeds another number by 7 and the product of the two numbers is 120. Find the number. I set it up like n+n+7=120, but the answers weren't right.

OpenStudy (bibby):

break up the statements. One number exceeds another number by 7 m = n+7 the product of the two numbers is 120. Find the number. m*n=120

OpenStudy (anonymous):

(n+7)n=120 n^2+7n=120 n^2+7-120=0 (n-8)(n+15) n=8 and n=-15?

OpenStudy (bibby):

-15*8 \(\ne \) 120

OpenStudy (bibby):

but 15 would work, eh

OpenStudy (anonymous):

yeah, did i work it out wrong then?

OpenStudy (bibby):

no, you worked out the quadratic correctly, I'm just unsure on why the answer came out like that

OpenStudy (anonymous):

haha weird!

OpenStudy (anonymous):

(x)*(x+7)=120 x=8

OpenStudy (anonymous):

one number is 8 and the other is 15

OpenStudy (anonymous):

depending on what number the question wants

OpenStudy (bibby):

oh ok, I see it now

OpenStudy (anonymous):

I see it too. Thank you!

OpenStudy (anonymous):

welcome good luck to you both :)